OpenStudy (anonymous):

write in slope intercept form: passes through (5,3) and (-5,-3)

OpenStudy (saifoo.khan):

C'mon, Good man knows every good things.

OpenStudy (anonymous):

hey!! i need help gosh, is that too much to ask???

OpenStudy (anonymous):

every goodman needs help sometime!!!!

OpenStudy (saifoo.khan):

lol, okay..

OpenStudy (saifoo.khan):

y = 3x/5

OpenStudy (anonymous):

thanks...ur arright!!

OpenStudy (anonymous):

slope is \[\frac{y{2} - y{1}}{x{2} - x{1}}\] and point-slope formula is \[y - y{1} = m(x - x{1})\] So the slope is 3/5 so use one of the points and the slope and you should get \[y = \frac{3}{5}x\]

OpenStudy (anonymous):

thanks i can under stand that better than ta oter guy!